• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 17:27:18 on June29,2010

    Entry number 329158

    This entry is a followup to: 329056

    keyword=Several BCMs unplugged and related checks

    I found that the input to the V2F in the middle counting room
    was unplugged for the following BCM signals: U10, U3, Unser,
    D3, and D1. I guess it's a casualty of PREX or HAPPEX.

    Meanwhile, D10 and U1 were plugged in. Looking at halog 329056,
    a short (2 min) run at 30 uA, the charge of 3.5 mC is about right
    in the halog for D10 and D1. So, at least two signals exist.

    To fix it, I plugged in the cables. I'm not 100% sure about
    Unser though; however, we probably don't care. Note, HAPPEX
    DAQ does not have several of these signals now. Also don't care.

    Fig 1 shows a selection of scaler channels during run 1336
    (30 uA on Pb). The T1 rate was 7 MHz. The canvas shows which
    BCMs work (see beam trip near event 250K) and some that do not.

    Good news: scalers are fine in the datastream and at least two
    current monitors work.

    On a related note, Stepan and I verified by plugging and unplugging
    cables that the scaler.map file is correct. So the students did
    this part correctly. I also verified, but using "fake" data,
    that the halla_scaler_process.tcl, which produces the halog end-run
    entries, is correct: it finds the right data and uses a correct
    scaler.map file. (These tests did NOT modify the halog.)


    FIGURE 1